OpenStudy (anonymous):

Evaluate sin^-1 1/2 . Round your answer to the nearest hundredth.

OpenStudy (jdoe0001):

the w0t? check your Unit Circle

OpenStudy (jdoe0001):

\(\bf sin^{-1}\left( \frac{1}{2} \right)\implies ?\) that is, "what is the angle whose sine is 1/2?"

OpenStudy (anonymous):

ok so we are going to start by drawing this on a coordinate plane

OpenStudy (anonymous):

since it is opposite over hypotenuse you label those 2 sides and you realize it is a special right triangle which will make the last side root 3

OpenStudy (anonymous):

we know that whatever angle is opposite the root 3 is 60 degrees, so that would make the inside angle we are looking for 30 degrees, which in radians is \[\frac{ \pi }{ 6}\]

OpenStudy (anonymous):

and that is your answer. any questions? do you understand now?

OpenStudy (anonymous):

how do i round to the nearest hundredth?????

OpenStudy (anonymous):

put pi/6 into your calculator and round it off

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

yep no problem!